# Climate impacts of equatorial superrotation

### v1

@@ 3.1 Atmospheric circulation
We first consider the response of the time-averaged atmospheric circulation to the imposed upper-tropospheric torque. Figure  1 shows the zonally-averaged zonal wind and the meridional mass streamfunction for both the control run ($F_{0}=0$ m.s11-1- $-1$, panel a) and the run with the largest prescribed torque ($F_{0}=4$ m.s11-1- $-1$, panel b).
Figure 1: Zonally- and time-averaged zonal wind (colors) and meridional mass streamfunction (contours with an interval of $10^{10}~{}\mathrm{kg.s}^{-1}$ and dashed lines indicating negative values), for experiments with $F_{0}=0$ m.s11-1- $-1$ (control run, panel a) and $F_{0}=4$ m.s11-1- $-1$ (panel b) and interactive SSTs.
@@
As the prescribed torque increases, we observe a transition to a state of superrotation (Fig. 1, panel b). This state is characterized by westerly winds over almost the entire atmosphere (except weak easterlies in the subtropics close to the surface), and a strong zonal jet of approximately $120$ m.s11-1- 1 at the equator, about twice the speed of the mid-latitude jets in the control run. The mean meridional circulation also undergoes major changes in its structure. The Hadley cell completely disappears, broken up into multiple smaller cells with weaker circulations; the mass meridional streamfunction decreases by an order of magnitude in the tropics. A residual Ferrell cell remains, albeit significantly weaker than in the control run and pushed poleward by about 20. Simulations for intermediate values of $F_{0}$, between $1$ and $3$ m.s11-1- $-1$, demonstrate that this transition is abrupt: the strength of the equatorial jet does not increase smoothly with the forcing amplitude but appears suddenly at $F_{0}=3$ m.s11-1- $-1$.
Figure 2: Zonally- and time-averaged zonal wind (colors) and meridional mass streamfunction (black contours, in units of $10^{10}~{}\mathrm{kg.s}^{-1}$), for $F_{0}=0$ m.s11-1- $-1$ (control run, panel a) and $F_{0}=4$ m.s11-1- $-1$ (panel b), for the fixed SST experiments.
Similarly, the zonally-averaged atmospheric circulations for the fixed-SST experiments are shown in Figure 2. The circulation in the control run (Fig. 2, panel a) is quite similar to that in the interactive-SST experiment. As expected, we observe a similar transition to equatorial superrotation (Fig. 2, panel b). However, the jet is weaker than in the interactive SST experiments, suggesting the existence of a positive feedback between surface temperature and the strength of the jet. This could be part of a self-sustaining mechanism for superrotation, should this positive feedback be strong enough to counterbalance damping in a setup with more typical forcings such as due to increases in CO2 concentrations or SST. The changes in the mean meridional circulation are also weaker, as relatively large circulation cells subsist at all latitudes. However, the direct cell in each hemisphere no longer resembles a Hadley cell, as it moves poleward by about 10 and weakens roughly by a factor of 3. In each hemisphere, a new indirect cell arises between the direct cell and the equator in the lower half of the atmosphere, while a weak direct cell subsists aloft. Like in the interactive-SST case, the Ferrell cells move poleward and decrease substantially in intensity.
@@ 3.2 Surface temperature
Figure 3 shows the zonally- and time-averaged surface temperature profiles for the control run ($F_{0}=0$ m.s-1.day-1) and in a forced superrotating state ($F_{0}=4$ m.s-1.day-1) for the simulations with a slab ocean. We first note that the control run is relatively warm compared to the real Earth system, particularly in the tropics, and has a strong equator-to-pole surface temperature gradient. These features of our idealized experiments are likely related to prescribing a relatively small, spatially uniform albedo and to the lack of a seasonal cycle for insolation, which leads to warmer tropics and colder polar regions. The globally-averaged surface temperature is nevertheless comparable to similar aquaplanet experiments (Hilgenbrink and Hartmann, 2018, e.g.).
In the superrotating state, surface temperature increases almost everywhere. The largest changes are in the mid-latitudes (between $30$ and $70^{\circ}$) and in the deep tropics (between 00 and $10^{\circ}$). In the subtropics (around $20^{\circ}$), the surface temperature changes induced by forced superrotation almost vanish. The globally-averaged surface temperature increases by $3~{}\mathrm{K}$, a temperature response comparable to that induced by a doubling of the atmospheric CO2 concentration (Forster et al., 2021). This result is remarkable: despite no external changes in the radiative properties of the atmosphere and without injecting any heat into the system, we nevertheless simulate a strong surface warming that can only be explained by the radiative response of the atmosphere to the imposed circulation change. These radiative forcings and feedbacks are explored in Section 4.
@@ 3.3 Precipitation
In the superrotating state, the globally-averaged precipitation rate increases slightly in the interactive-SST case (Fig. 4, a), but less than the $2-3$.K11-1- 1 usually found in global warming experiments (O’Gorman et al., 2012). In the fixed-SST case, however, the globally-averaged precipitation rate decreases slightly: these two results are in agreement with tropospheric energy budget arguments, based on the approximate balance between radiative cooling and latent heat release above cloud base (O’Gorman et al., 2012). Perhaps more interestingly, we find a drastic change in the meridional distribution of precipitation in the superrotating state: Precipitation is strongly decreased in the tropics, slightly decreased in the mid-latitudes (the two regions where the precipitation rate has local maxima in the control climate) and increased in the subtropics and at high latitudes, leading overall to a much more homogeneous spatial distribution of precipitation compared to the current climate. These changes are qualitatively similar between the interactive- and fixed-SST experiments, although there are some differences: In the superrotating states, the emergent precipitation maxima in the subtropics are further away from the equator in the fixed-SST runs, and there are no maxima of precipitation at $30^{\circ}$ of latitude as in the interactive-SST runs. These changes in the precipitation profiles are consistent with the circulation changes described above. Indeed in the current climate, the Hadley cells transport moisture from the subtropics towards the equator (Hartmann, 2016). The weakening of this overturning circulation, shown in Figure 1, provides a possible explanation for the changes in tropical precipitation. The processes related to changes in the precipitation distribution are discussed in more detail in Section 6
@@ 4 From circulation changes to energy budget changes
In order to link the climate changes described in the previous section to the imposed circulation change, we will investigate how these circulation changes influence the distribution of energy in the climate system. In particular, we discuss the roles of the mean meridional circulation and of the eddies in Section 4.4.1, and then use the methods introduced in Section 2 to discuss the radiative balance changes in Section 4.4.2 and Section 4.4.3.
@@ 4.1 Meridional energy transport
After applying a correction to the meridional wind in order to compensate the numerical errors in the closure of the energy balance (Bangalath and Pauluis, 2020), we compute explicitly the meridional MSE flux and decompose it into mean meridional circulation (MMC) and eddy components (Fig. 5). Here, the eddies include both transient and standing eddies, but since the boundary conditions are axisymmetric in our problem, standing eddies are not expected to play a very important part.
Figure 5: a) total atmospheric meridional energy transport, positive northward, along with its mean meridional circulation and eddy components in the control run with interactive SST. Bottom: Transport changes between the superrotating ($F_{0}=4$ m.s11-1- $-1$) and control ($F_{0}=0$ m.s11-1- $-1$) runs, for interactive (b) and fixed SST (c).
Figure 6: Changes in meridional energy transport divergence between the superrotating run ($F_{0}=4$ m.s11-1- $-1$) and the control run ($F_{0}=0$ m.s11-1- $-1$), averaged over three regions: tropics ($15^{\circ}$ to $15^{\circ}$), subtropics ($40^{\circ}$ to $15^{\circ}$ and $15^{\circ}$ to $40^{\circ}$), and mid-/high-latitudes ($90^{\circ}$ to $40^{\circ}$ and $40^{\circ}$ to $90^{\circ}$), for interactive- (blue bars) and fixed-SST runs (pink bars). The mean meridional circulation (orange arrows) and eddies (green arrows) components of these changes are shown.
@@
In the superrotating state (Fig. 5, panels b and c), the most evident change is that the poleward eddy energy flux (green curve) increases at all latitudes, while the magnitude of the energy flux associated with the MMC (orange curve) decreases at all latitudes. The former is broadly consistent with the appearance of a turbulent jet at the equator, while the latter can be explained qualitatively by the collapse of the Hadley cells at low latitudes and the poleward shift of the Ferrell cells at high latitudes (the two regions are separated by a buffer region where energy transport changes due to the MMC are small, around 40).
The partial compensation of these opposing effects determines the pattern of total poleward energy transport changes (blue curve), which is negative in the tropics (roughly until $20^{\circ}$ for the interactive-SST runs and $15^{\circ}$ for the fixed-SST runs), with a minimum near $10^{\circ}$, positive in the mid-latitudes, until about $55^{\circ}$ ($50^{\circ}$ in the fixed-SST runs) and with a maximum near $35^{\circ}$ (30 in the fixed SST runs), and negative further poleward.
This indicates that the MMC drives the total energy transport changes in the tropics and high latitudes, and the eddies do in the subtropics. As a result, the change in net MSE flux divergence is negative in the tropics and mid-latitudes, and positive in the subtropics (Fig. 6).
@@
4.2 Radiative budget changes and their relation to temperature and humidity
@@
Figure 7: TOA flux changes (panel a) across the transition to superrotation (from $F_{0}=0$ to $F_{0}=4$ m.s11-1- $-1$), and the decomposition of these changes into radiative forcings (panel b) and feedbacks (panel c) computed with radiative kernel. Panel a corresponds to the interactive-SST runs, panel b to the fixed-SST simulations and panel c is the difference between the two. Black curves represent the radiative budgets computed by the radiation scheme in the model, blue curves are the budgets reconstructed using the radiative kernels, which can be broken down into the Planck response (yellow curves), specific humidity response (red curves) and lapse-rate response (green curves). Numerical values given in the legend correspond to the global averages of the various components.
@@
We show in Figure 7 the TOA radiative budget changes (black line) in the interactive-SST runs (panel a), in the fixed-SST runs (the radiative forcing, panel b) and the difference between the two (i.e., the feedbacks, panel c). The meridional structure of the TOA radiative changes in the interactive-SST runs is, as expected, in agreement with the analysis of the energy transport changes: the excess of energy transport in the deep tropics (between $10^{\circ}$ and $10^{\circ}$) and in the mid- and high-latitudes (poleward of $40^{\circ}$) is balanced by increased OLR, while the opposite holds in a broad subtropical region (between $10^{\circ}$ and $40^{\circ}$). The net TOA radiative budget changes (panel a) are dominated by the radiative forcing (panel b) in the low-latitudes (between $40^{\circ}$ and $40^{\circ}$) and by radiative feedbacks at higher latitudes (poleward of $40^{\circ}$).
@@
To relate these changes in the TOA radiative budget to physical fields, we use radiative kernels (as described in Section 22.3). First, we compare the total radiative budget changes computed using the kernels (Fig. 7, blue curves) with the changes computed directly from the model outputs (Fig. 7, black curves). Overall, the meridional structures of the two curves are very similar, though there are differences due both to the linearization inherent to the kernel method and to the fact that we use kernels constructed with a different radiation scheme. In particular, the domination of the forcing and feedback contributions at low and mid-/high-latitudes respectively is even more pronounced in the radiative kernel budget. In the following, we will embrace this approximation and explain qualitatively why the low latitudes are dominated by the forcing and the mid- and high-latitudes by the feedbacks. We note that the smaller contributions of the forcing at mid- and high-latitudes is at odds with the larger energy transport changes reported in those regions in the fixed-SST runs in Section 4.4.1. This can be explained by the fact that the surface fluxes, as noted above, are not constrained in this case, and indeed their changes absorb most of the energy transport changes, leaving the TOA radiative budget largely unaltered.
We proceed with the radiative kernel approach to decompose the forcing, feedbacks, and their sum (as introduced in Section 22.2) into contributions from changes in specific humidity (Fig. 7, red curves), lapse rate (Fig. 7, green curves) and surface temperature (the Planck response, yellow curve). There is no region where a single component clearly dominates, but we can nevertheless relate each contribution to changes in the associated physical fields.
We start with the meridional structure of the radiative forcing (i.e., the TOA anomaly induced in the fixed-SST runs by prescribing $F_{0}=4$ m.s11-1- $-1$). We show in Figure 8 (panel b) and Figure 9 (panel a) the time- and zonally-averaged specific humidity and temperature changes, respectively. In the superrotating state, the atmosphere is drier than the control in the tropics, moister in the mid-latitudes, and slightly drier close to the poles. This explains qualitatively the meridional structure of the contribution of specific humidity changes to the radiative forcing (Fig. 7, panel b): drier air results in decreased optical depth, increased OLR, and negative forcing, all else equal (conversely for moister air). Similarly, Figure 9 (panel b) shows that the lapse rate increases in the tropics, decreases in mid-latitudes, and increases again close to the poles, which matches the meridional structure of the lapse rate contribution to radiative forcing (Fig. 7, panel b).
Figure 8: Time- and zonally-averaged fractional specific humidity changes between the $F_{0}~{}=~{}4$ m.s11-1- $-1$ and $F_{0}~{}=~{}0$ m.s11-1- $-1$ experiments with interactive (panel a) and fixed SST (panel b). Here and for Figure 9, the dashed and solid black lines show the time- and zonally-averaged emission heights for $F_{0}=0$ and $F_{0}=4$ m.s11-1- $-1$, respectively.
Figure 9: Time- and zonally-averaged atmospheric temperature changes between the $F_{0}~{}=~{}4$ m.s11-1- $-1$ and $F_{0}~{}=~{}0$ m.s11-1- $-1$ experiments with interactive (panel a) and fixed SST (panel b).
However, the way the two effects combine into the net radiative forcing (Fig. 7, panel b) is more subtle. To understand it, we consider jointly the emission level changes due to specific humidity changes (Figs. 8 and  9) and the spatial distribution of temperature changes. Close to the equator, the drier atmosphere results in a significantly lower emission level, which is not compensated for by the lapse rate changes. Although the temperature in the superrotating state at the new emission level is cooler than it was at the same level in the control run, it is warmer than the temperature at the emission level in the control run, hence the OLR is larger: specific humidity effects dominate. Moving poleward, the competition between the two effects progressively changes: while the atmosphere is still drier compared to the superrotating state, it is less so and the emission level changes are smaller. At the same time, the lapse rate changes amplify. The combined change results in decreased OLR changes, until it vanishes around 15 and even becomes negative when lapse rate effects dominate between $15^{\circ}$ and $20^{\circ}$. Between $20^{\circ}$ and $40^{\circ}$, the two effects revert (Fig. 9): the lapse rate changes decrease until they become negative around $30^{\circ}$ and poleward, and the atmosphere progressively becomes moister than the control run (Fig. 8), resulting in higher emission levels. As a result, the lapse rate contribution to the forcing decreases with latitude in this region (Fig. 7, panel b) while the specific humidity contribution increases before levelling off. The fact that the OLR is smaller than in the control run in this region (the blue curve is positive) is therefore driven first by the specific humidity effects in the first half of the region, then by lapse rate effects in the second half. Going further poleward, the two effects essentially cancel out, resulting in a vanishing radiative forcing up to very high latitudes.
For the interactive-SST experiments, Figure 8 (panel a) shows that the specific humidity changes are smaller than in the fixed-SST case equatorward of $20^{\circ}$ (in this region, the emission level remains the same) — the superrotating state is actually moister except in the region where the OLR value is larger. Lapse rate changes are essentially smaller everywhere (Fig. 9, panel a). This explains the structure of the radiative feedback curves (Fig. 7, panel c), whose sign is opposite to the radiative forcing everywhere for the lapse rate component (green curve), and opposite in the tropics but same sign in the mid-latitudes for the specific humidity component (red curve; in other words specific humidity acts as a positive feedback everywhere). As a result, the specific humidity effect on the net radiative budget changes (Fig. 7, panel a) vanishes in the tropics while the lapse rate component remains relatively small everywhere. The total radiative budget changes are dominated by the Planck response in the deep tropics (with a minor contribution from the lapse rate), and become dominated by the imperfect cancellation between the Planck and specific humidity responses at higher latitudes.
Overall, the TOA radiative budget changes are consistent with changes in physical quantities in the atmosphere, such as specific humidity and temperature, but result from cancellations between the effects associated with these variables. In the following section, we suggest an alternative decomposition of the radiative budget changes, which offers a clearer interpretation.
4.3 Radiative budget changes and their relation to relative humidity
Given the strong correlation between the effects of specific humidity and lapse rate changes on the TOA budget (see Section 44.2), it is natural to try to interpret these changes using the alternative decomposition introduced by Held and Shell (2012) and Ingram (2012) (see also Jeevanjee et al. (2021)), which uses relative humidity instead of specific humidity as a state variable. This means that for the two other terms, namely the Planck and lapse rate components, it is relative humidity that is considered constant rather than specific humidity. This helps to avoid non-physical results, such as a temperature decrease with fixed specific humidity leading to supersaturated air, and it also helps disentangle the two competing effects of specific humidity and lapse rates on radiative fluxes. The radiative kernel for this decomposition can be computed from the original decomposition, as follows:
$\begin{pmatrix}\delta R_{q}\\ \delta R_{\mathrm{P}}\\ \delta R_{\Gamma}\end{pmatrix}=\begin{pmatrix}K^{\mathrm{wv}}\frac{\delta\log(q)}{\partial\log(q)/\partial T}\\ \left(K^{T_{\mathrm{s}}}+K^{T}\right)\delta T_{\mathrm{s}}\\ K^{T}\left(\delta T-\delta T_{\mathrm{s}}\right)\end{pmatrix}\xrightarrow[\mathrm{{state\,variables}}]{\mathrm{Change\,of}}\begin{pmatrix}\widetilde{\delta R}_{\mathrm{RH}}\\ \widetilde{\delta R}_{\mathrm{P}}\\ \widetilde{\delta R}_{\Gamma}\end{pmatrix}=\begin{pmatrix}K^{\mathrm{wv}}\frac{\delta\log\left(\mathrm{RH}\right)}{\partial\log(q)/\partial T}\\ \left(K^{T_{\mathrm{s}}}+K^{T}+K^{\mathrm{wv}}\right)\delta T_{\mathrm{s}}\\ \left(K^{T}+K^{\mathrm{wv}}\right)\left(\delta T-\delta T_{\mathrm{s}}\right)\end{pmatrix},$ (5)
where $\delta R_{\left[q,\mathrm{P},\Gamma\right]}$ and $\widetilde{\delta R}_{\left[\mathrm{RH},\mathrm{P},\Gamma\right]}$ are the radiative responses associated to the different components from the original decomposition and the alternative decomposition using relative humidity, respectively. $K^{\left[\mathrm{wv},T_{\mathrm{s}},T\right]}$ are the radiative kernels, respectively, associated with specific humidity, surface temperature, and air temperature. In practice, we compute the relative humidity response by writing it as $\widetilde{\delta R}_{\mathrm{RH}}=K^{\mathrm{wv}}\frac{\delta\log(q)}{\partial\log(q)/\partial T}-K^{\mathrm{wv}}\delta T$. This relation can be obtained in a straightforward manner, by writing $\delta R$ in terms of partial derivatives.
@@ 4.3 Radiative budget changes and their relation to relative humidity
Figure 10: Total response and the different components of the alternative decomposition of TOA radiative anomalies using relative humidity, computed using radiative kernels, between the $F_{0}=0$ and $F_{0}=4$ m.s11-1- $-1$ experiments. Panel a corresponds to the interactive-SST runs, panel b to the fixed-SST runs, and panel c is the difference between the two. Numerical values given in the legend correspond to the global average of the associated component.
@@
We show in Figure 10 this alternative decomposition of the TOA flux changes in the interactive-SST runs (panel a), in the fixed-SST runs (i.e., the radiative forcing, panel b) and the difference (i.e., the feedbacks, panel c). The results are strikingly clear: the lapse rate contributions are now vanishingly small for both the forcing and feedbacks (and, therefore, for the net radiative changes also). This indicates that the lapse rate effects described in the previous section were almost entirely due to changes in the moist adiabatic lapse rate, i.e. associated with Clausius-Clapeyron effects. It follows that the radiative forcing (panel b) is now almost entirely explained by the relative humidity changes (red curve), which is confined to the tropics and subtropics and largely vanishes elsewhere. We can therefore compare the meridional structure of the radiative forcing to the relative humidity changes, shown in Figure 11 (panel b for the fixed-SST runs).
@@
Figure 11: Time- and zonally-averaged fractional relative humidity changes between the $F_{0}~{}=~{}4$ m.s11-1- $-1$ and $F_{0}~{}=~{}0$ m.s11-1- $-1$ experiments, for the interactive-SST (panel a) and fixed-SST (panel b) experiments.
@@
Indeed, the OLR increase in the tropics (within $15^{\circ}$ of the equator) corresponds to a decrease in relative humidity, while the OLR decrease in the subtropics and until $40^{\circ}$ is associated with an increase in relative humidity. Note that although the relative humidity changes are stronger in the subtropics, the two effects lead to radiative forcings of similar magnitudes (but opposite in sign) because the radiative kernel is larger closer to the equator. For the same reason, the relative humidity changes poleward of $40^{\circ}$ do not lead to appreciable radiative forcing.

### v3

@@ 3.1 Atmospheric circulation
We first consider the response of the time-averaged atmospheric circulation to the imposed upper-tropospheric torque. Figure 1 shows the zonally-averaged zonal wind and the meridional mass streamfunction for both the control run ($F_{0}=0$ m.s1-1.day1-1, panel a) and the run with the largest prescribed torque ($F_{0}=4$ m.s1-1.day1-1, panel b).
Figure 1: Zonally- and time-averaged zonal wind (colors) and meridional mass streamfunction (contours with an interval of $10^{10}~\mathrm{kg.s}^{-1}$ and dashed lines indicating negative values), for experiments with $F_{0}=0$ m.s1-1.day1-1 (control run, panel a) and $F_{0}=4$ m.s1-1.day1-1 (panel b) and interactive SSTs.
@@
As the prescribed torque increases, we observe a transition to a state of superrotation (Fig. 1, panel b). This state is characterized by westerly winds over almost the entire atmosphere (except weak easterlies in the subtropics close to the surface), and a strong zonal jet of approximately $150$ m.s1-1 at the equator, about twice the speed of the mid-latitude jets in the control run. The mean meridional circulation also undergoes major changes in its structure. The Hadley cell is strongly weakened; the mass meridional streamfunction decreases by a factor $3$ in the tropics. A residual Ferrell cell remains, albeit significantly weaker than in the control run and shifted poleward by about 20. Simulations for intermediate values of $F_{0}$, between $1$ and 33 m.s1-1.day1-1 (not shown), demonstrate that this transition is abrupt: the strength of the equatorial jet does not increase smoothly with the forcing amplitude. Instead, it appears suddenly at $F_{0}=3$ m.s1-1.day1-1.
Figure 2: Zonally- and time-averaged zonal wind (colors) and meridional mass streamfunction (black contours, in units of $10^{10}~\mathrm{kg.s}^{-1}$), for $F_{0}=0$ m.s1-1.day1-1 (control run, panel a) and $F_{0}=4$ m.s1-1.day1-1 (panel b), for the fixed SST experiments.
Similarly, the zonally-averaged atmospheric circulations for the fixed-SST experiments are shown in Figure 2. The circulation in the control run (Fig. 2, panel a) is quite similar to that in the interactive-SST experiment. As expected, we observe a similar transition to equatorial superrotation (Fig. 2, panel b). However, the jet is weaker than in the interactive SST experiments, suggesting the existence of a positive feedback between surface temperature and the strength of the jet. This could be part of a self-sustaining mechanism for superrotation, should this positive feedback be strong enough to counterbalance damping in a setup with more typical forcings such as due to increases in CO2 concentrations or SST. The changes in the mean meridional circulation are also weaker but follow the same structure as in the interactive SST case: the intensities of both the direct and indirect cells decrease and they are shifted poleward.
@@ 3.2 Surface temperature
Figure 3 shows the zonally- and time-averaged surface temperature profiles for the control run ($F_{0}=0$ m.s-1.day-1) and in a forced superrotating state ($F_{0}=4$ m.s-1.day-1) for the simulations with a slab ocean. We first note that the control run is relatively warm compared to the real Earth system, particularly in the tropics, and has a strong equator-to-pole surface temperature gradient. These features of our idealized experiments are likely related to prescribing a relatively small, spatially uniform albedo and to the lack of a seasonal cycle for insolation, which leads to warmer tropics and colder polar regions. The globally-averaged surface temperature is nevertheless comparable to similar aquaplanet experiments (hilgenbrinkResponseHadleyCirculation2018, e.g.).
In the superrotating state, surface temperature increases almost everywhere. The largest changes are in the mid-latitudes (between $30$ and $70^{\circ}$) and in the deep tropics (between 0 and $10^{\circ}$). In the subtropics (around $20^{\circ}$), the surface temperature changes induced by forced superrotation almost vanish. The globally-averaged surface temperature increases by approximately $3.5~\mathrm{K}$, a temperature response comparable to that induced by a doubling of the atmospheric CO2 concentration (IPCC_2021_WGI_Ch_7). This result is remarkable: despite no external changes in the radiative properties of the atmosphere and without injecting any heat into the system, we nevertheless simulate a strong surface warming that can only be explained by the radiative response of the atmosphere to the imposed circulation change. These radiative forcings and feedbacks are explored in Section 4.
@@ 3.3 Precipitation
In the superrotating state, the globally-averaged precipitation rate increases slightly in the interactive-SST case (Fig. 4, a), but less than the 232-3 %.K1-1 usually found in global warming experiments (ogormanEnergeticConstraintsPrecipitation2012). In the fixed-SST case, the increase in globally-averaged precipitation rate is even weaker: these two results are in agreement with tropospheric energy budget arguments, based on the approximate balance between radiative cooling and latent heat release above cloud base (ogormanEnergeticConstraintsPrecipitation2012). Perhaps more interestingly, we find a drastic change in the meridional distribution of precipitation in the superrotating state: Precipitation is strongly decreased in the tropics, slightly decreased in the mid-latitudes (the two regions where the precipitation rate has local maxima in the control climate) and increased in the subtropics and at high latitudes, leading overall to a much more homogeneous spatial distribution of precipitation compared to the current climate. These changes are qualitatively similar between the interactive- and fixed-SST experiments, although there are some differences: In the superrotating states, the emergent precipitation maxima in the subtropics are further away from the equator in the fixed-SST runs, and there are no maxima of precipitation at $30^{\circ}$ latitude as in the interactive-SST runs. These changes in the precipitation profiles are consistent with the circulation changes described above. Indeed in the current climate, the Hadley cells transport moisture from the subtropics towards the equator (hartmannGlobalPhysicalClimatology2016). The weakening of this overturning circulation, shown in Figure 1, provides a possible explanation for the changes in tropical precipitation. The processes related to changes in the precipitation distribution are discussed in more detail in Section 6
@@ 4 From circulation changes to energy budget changes
In order to link the climate changes described in the previous section to the imposed circulation change, we will investigate how these circulation changes influence the distribution of energy in the climate system. In particular, we discuss the roles of the mean meridional circulation and of the eddies in Section 4.4.1, and then use the methods introduced in Section 2 to discuss the radiative budget changes in Section 4.4.2.
@@ 4.1 Meridional energy transport
After applying a correction to the meridional wind in order to compensate the numerical errors in the closure of the energy balance (bangalathNewMassFlux2020), we compute explicitly the meridional MSE flux and decompose it into mean meridional circulation (MMC) and eddy components (Fig. 5). Here, the eddies include both transient and standing eddies, but since the boundary conditions are axisymmetric in our problem, standing eddies are not expected to play a very important part.
Figure 5: a) total atmospheric meridional energy transport, positive northward, along with its mean meridional circulation and eddy components in the control run with interactive SST. Bottom: Transport changes between the superrotating ($F_{0}=4$ m.s1-1.day1-1) and control ($F_{0}=0$ m.s1-1.day1-1) runs, for interactive (b) and fixed SST (c).
Figure 6: Changes in meridional energy transport divergence between the superrotating run ($F_{0}=4$ m.s1-1.day1-1) and the control run ($F_{0}=0$ m.s1-1.day1-1), averaged over three regions: tropics ($12^{\circ}$S to $12^{\circ}$N), subtropics ($40^{\circ}$S to $12^{\circ}$S and $12^{\circ}$N to $40^{\circ}$N), and mid-/high-latitudes ($90^{\circ}$S to $40^{\circ}$S and $40^{\circ}$N to $90^{\circ}$N), for interactive- (top panel) and fixed-SST runs (bottom panel). The total (blue bars), mean meridional circulation (orange bars) and eddies (red bars) components of these changes are shown.
@@
In the superrotating state (Fig. 5, panels b and c), the most evident change is that the poleward eddy energy flux (green curve) increases at all latitudes, while the magnitude of the energy flux associated with the MMC (orange curve) mostly decreases in the tropics and subtropics, and becomes weakly positive in the mid-latitudes. The former is broadly consistent with the appearance of a turbulent jet at the equator, while the latter can be explained qualitatively by the collapse of the Hadley cells at low latitudes and the poleward shift of the Ferrell cells at high latitudes (the two regions are separated by a buffer region where energy transport changes due to the MMC are small, around 40).
The partial compensation of these opposing effects determines the pattern of total poleward energy transport change (blue curve): in the interactive SST case, it is negative in the tropics (roughly until $12^{\circ}$ latitude), with a minimum near $10^{\circ}$, and positive at higher latitudes with a maximum near $40^{\circ}$ latitude. This indicates that the MMC mostly drives the total energy transport changes in the tropics and high latitudes, and the eddies do in the subtropics. As a result, the change in net MSE flux divergence is negative in the tropics and mid-latitudes, and positive in the subtropics (Fig. 6).
@@
4.2 Radiative budget changes and their relation to relative humidity
@@
We show in Figure 7 the TOA radiative budget changes (black line) in the interactive-SST runs (panel a), in the fixed-SST runs (the radiative forcing, panel b) and the difference between the two (i.e., the feedbacks, panel c). The meridional structure of the TOA radiative changes in the interactive-SST runs is, as expected, in agreement with the analysis of the energy transport changes: the excess of energy transport in the deep tropics (between $10^{\circ}$S and $10^{\circ}$N) and in the mid- and high-latitudes (poleward of $40^{\circ}$) is balanced by increased OLR, while the opposite holds in a broad subtropical region (between $10^{\circ}$ and $40^{\circ}$). The net TOA radiative budget changes (panel a) are dominated by the radiative forcing (panel b) in the low-latitudes (between $40^{\circ}$S and $40^{\circ}$N) and by radiative feedbacks at higher latitudes (poleward of $40^{\circ}$).
To relate the changes in TOA radiative budget, in particular the radiative forcing and feedback components, to physical fields, we use radiative kernels (as described in Section 22.3). We choose the relative humidity-based decomposition introduced by Held2012a and Ingram2012 [see also Jeevanjee2021], in which relative humidity is held fixed in the Planck and lapse-rate terms. In addition to the general arguments for this method, including that it avoids unphysical scenarios (e.g. cooling at fixed specific humidity leading to supersaturation) and disentangles the humidity and lapse rate effects related to the Clausius-Clapeyron relation, in our particular case it provides a much simpler interpretation of the radiative budget changes, as discussed in Appendix A. This relative humidity-based decomposition results in contributions from changes in relative humidity (Fig. 7, pink curves), lapse rate (orange curves), and surface temperature (Planck response, blue curve) to changes in TOA radiation. Similar decompositions apply for radiative forcing and feedbacks: $F_{\mathrm{tot,kernel}}=F_{\Gamma}+F_{\mathrm{RH}}$ and $(\lambda\delta T_{\mathrm{s}})_{\mathrm{tot,kernel}}=-{(\lambda\delta T_{\mathrm{s}})}_{\mathrm{T_{\mathrm{s}}}}-{(\lambda\delta T_{\mathrm{s}})}_{\Gamma}-{(\lambda\delta T_{\mathrm{s}})}_{\mathrm{RH}}$; these components are represented in Fig. 7 with the same color code. All these terms can be expressed using the original kernel computed in liuRadiativeKernelsIsca2020, as explained in Appendix B.
@@
First, we compare the total radiative budget changes computed using the kernels (Fig. 7, dashed black curves) with the changes computed directly from the model outputs (Fig. 7, solid black curves). Overall, the meridional structures of the two curves are very similar, though there are differences due both to the linearization inherent to the kernel method and to the fact that we use kernels constructed with a different radiation scheme. In particular, the domination of the radiative forcing and feedback contributions at low (below $40^{\circ}$) and mid-/high-latitudes (above $40^{\circ}$) respectively is even more pronounced in the radiative kernel budget. In the following, we will embrace this approximation and explain qualitatively why the low latitudes are dominated by the forcing and the mid- and high-latitudes by the feedbacks. We note that the smaller contributions of the radiative forcing at mid- and high-latitudes is at odds with the larger energy transport changes reported in those regions in the fixed-SST runs in Section 4.4.1. This can be explained by the fact that the surface fluxes, as noted above, are not constrained in this case, and indeed their changes absorb most of the energy transport changes, leaving the TOA radiative budget largely unaltered.
@@
Figure 7: Total response and the different components of the decomposition of TOA radiative anomalies using relative humidity, computed using radiative kernels, between the $F_{0}=0$ and $F_{0}=4$ m.s1-1.day1-1 experiments. Panel a corresponds to the interactive-SST runs, panel b to the fixed-SST runs, and panel c is the difference between the two. Numerical values given in the legend correspond to the global average of the associated component.
@@
We now consider the decomposition of the TOA radiative flux changes in the interactive-SST runs (panel a), in the fixed-SST runs (i.e., the radiative forcing, panel b) and the difference (i.e., the feedbacks, panel c) into their lapse rate, relative humidity and Planck components (Figure 7, color curves). The results are strikingly clear: the lapse rate contributions are vanishingly small for both the radiative forcing and feedbacks (and, therefore, for the net radiative changes also), and the radiative forcing (panel b) is almost entirely explained by the relative humidity changes (red curve), which is confined to the tropics and subtropics and largely vanishes elsewhere. We can therefore compare the meridional structure of the radiative forcing to the relative humidity changes, shown in Figure 8 (panel b for the fixed-SST runs).
@@
Figure 8: Time- and zonally-averaged fractional relative humidity changes between the $F_{0}~=~4$ m.s1-1.day1-1 and $F_{0}~=~0$ m.s1-1.day1-1 experiments, for the interactive-SST (panel a) and fixed-SST (panel b) experiments.
@@
Indeed, the OLR increase in the tropics (within $12^{\circ}$ latitude of the equator) corresponds to a decrease in relative humidity, while the OLR decrease in the subtropics and until $40^{\circ}$ is associated with an increase in relative humidity. Note that although the relative humidity changes are stronger in the subtropics, the two effects lead to radiative forcings of similar magnitudes (but opposite in sign) because the radiative kernel is larger closer to the equator. For the same reason, the relative humidity changes poleward of $40^{\circ}$ do not lead to appreciable radiative forcing.